1. 0.0050 grams of a chemical is dissolved in 75 grams of water. What is the concentration of the chemical in parts per million? 




1
Expert's answer
2022-04-04T17:02:02-0400

solute = 0.0050g

solvent = 75g


concentration in ppm = "\dfrac{solute}{(solute + soluvent)}X 1 000 000"


concentration in ppm = "\dfrac{0.0050g}{(0.0050g + 75g)}" x 1000000 = 66.66ppm